Haruki has $2000 that he wishes to invest for the next two years. One year GICs are currently paying 8% while two year GICs are paying 12 % compounded annually. Economists are predicting that interest rates will rise by the end of he year. what is the minimum interest rate Haruki would need in year two, to make the one year GIC better than the two year GIC ?(Show your calculations)
Answer #1

P=2000,r1=8%,

F=P(1+r)^2

=2000(1+0.12)^2

2508.8

hence

2508.8=2000(1+r)

1.25=1+r

r=25%, hence 25% is required to make one year better than two.